Information Technology Staff Meeting Notes
December 12, 2008
December 12, 2008
Present: Thomas Bielicke, Cynthia Bowker, Laurie Brennan, Will Chaney, Logan Clapp, Matthew Daley, Debra Fetzer, Jason Fleming, Winston Harris, Michael Jay, Paul McDonough, Bobbie Meng, Tom Minton, Chris Nicolich, Amy Polk, Patricia Ruwell, Evan Wack, Stephen WilliamsAnnouncements
Ongoing IT Updates
- Stephen finished his tests. He still has an overdue project to complete.
- Paul took his last undergrad test and has been cleared for graduation.
- IE7 - Microsoft advises to set the security to high. It can be done with Group Policy. If it is acceptable, we will do it in Group Policy. There are three sites that legitimately require IE7: PeopleFirst, PeopleSoft and Mister Paper.
Workstation Support Unit
- Exchange Failure Status - most problems have been fixed. Backup Exec - we've been able to restore current backups. We are still missing the data from November 12th -18th.
- Exchange and AD - will test on the IT department first - maybe next week. After the holiday break, if the funding is approved, we will move the remainder of the staff.
- No news about Tivoli - we are waiting for Director approval.
- SAN migration this weekend - Admin and Projects.
- XP SP3 fix - will be done over the holidays.
- Inventory - we'll do it with laptops and update Grover as we go. We have decided to wait until after the holidays - and AD.
- Public updates - high ends - Steady State lengthens the reboot.
Software Projects Status
- Aleph update ran on Sunday - required printer changes. Law is still having problems.
- FCLA is doing a Tivoli backup of ALEPH reporting - no major changes should be made from Dec. 24 - Jan 3.
- CSUL - voted to go to Aleph 19. Will test in the spring and migrate in the summer.
- Mark - JPEG2K backup - wants more frequent backups. Cynthia will forward the email to Will.
- New version of Grover was released today. Hardware reports are now on the Hardware page. Inventory can display in Excel. Physical location added. VOIP check box added.
- GatorScholar - 3rd level domain request - no response.
- ARL stats - ongoing.
- Michael is working on another project for Resource Services - CatQC. Involves flipping call numbers and logging occurrences.
- DLC Tracking database - problem this week - Winston undid a change and it was fixed.
- Greenstone - problems this AM - updates installed. It is working now. Winston had to restart the SQL server part of UFDC.
Students and Department
- The new homepage design is in its 25th revision. The Dean now wants it as soon as possible. Tom is meeting with Michele Crump on the timeline. Will beta test in the spring and roll out by the summer.
- Tom will be out on sabbatical for the fall semester. Matthew may also be out on paternity leave about that time.
- The new home page will also require redesigning all top level pages. We are working on a new template.
- IRB is interested in Web survey. They may require a consent form. Survey Monkey is part of the issue.
IT Security Management
- Finals week is next week. We should have coverage the first half of the week, and then they will be dropping off.
- Paul Losch inquired about the old Latin American tapes we have (now stored on VTS at CNS). He had an inquiry from ICPSR in Michigan. They are interested in archiving the material. Debra gave Paul a list of the volsers. There are about 125. It's possible that we may be asked to FTP the data to their server. Paul is informing them of our holdings and will notify us when any further action is necessary.
- Meeting held.